how many moles of sodium sulfate will be produced from 25.5 g

 of sodium hydroxide?



1
Expert's answer
2022-03-23T02:44:03-0400

We know that we get sodium sulphate from sodium hydroxide when sodium hydroxide reacts with sulphuric acid reaction can be shown as follows

"2NaOH+ H_2SO_4 \longrightarrow Na_2SO_4 + 2H_2O"


From above reaction we can find that 2 mole of sodium hydroxide will give 1 mole of sodium sulphate


mass of sodium hydroxide given = 25.5


mole of sodium hydroxide = "\frac{mass}{molar mass}= \frac{25.5}{40}=0.6375"


mole of Na2SO4 produce = "0.5 \times 0.6375 = 0.31875" mole


mass of sodium sulphate produce = "mole \times molar mass= 0.31875 \times 142= 45.26"


mass of sodium sulphate produce = 45.26 g